Find how many 4 digits number are even and how many 4 digits number are odd.
Answer by sudhanshu_kmr(1152)


total no. of 4 digit even number = 9 * 10 * 10 * 5 = 4500
(here, at first place only 1-9, at last any even number and other places any digit 0-9)


total no. of 4 digit odd number = 9 *10*10* 5 = 4500
(similar to first condition ,at first place only 1-9, at last any odd numbers and other places any digit 0-9)